# SPECIAL RELATIVITY

### v11

@@ 21 Appendix
[…]
×$\times\left[\left({v_{x}(z^{\prime})\over{c}}-(\theta_{x}-\eta_{x})\right){\vec{e_{x}}}+\left({v_{y}(z^{\prime})\over{c}}-(\theta_{y}-\eta_{y})\right){\vec{e_{y}}}\right]~{},$ (160)
@@
[…]
@@
[…]
×$\times\frac{1}{z-z^{\prime}}{\rho_{\bot}}\left(\vec{r^{\prime}}_{\bot}\right)\left(\frac{\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}}{z-z^{\prime}}\right)~{}.$
@@
×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\int_{0}^{\infty}\frac{d\xi}{\xi}\exp\left[+i\omega\frac{\mid\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}\mid^{2}}{2c\xi}-\frac{i\omega\xi}{2c\gamma^{2}}\right]\Bigg{\}}$
@@
[…]
@@
[…]

### v12

@@ 22 Appendix
$\vec{\widetilde{{E}}}(z_{0},\vec{r}_{0},\omega)=-{i\omega e\over{c^{2}}z_{0}}\int_{-\infty}^{\infty}dz^{\prime}{\exp{\left[i\Phi_{T}\right]}}$ (159)
$\times\left[\left({v_{x}(z^{\prime})\over{c}}-(\theta_{x}-\eta_{x})\right){\vec{e_{x}}}+\left({v_{y}(z^{\prime})\over{c}}-(\theta_{y}-\eta_{y})\right){\vec{e_{y}}}\right]~,$ (160)
@@
$\vec{{\bar{E}}}(z,\vec{r}_{\bot},\omega)$ $=$ $-4\pi\int_{-\infty}^{\infty}dz^{\prime}\int d\vec{r^{\prime}}_{\bot}\left(\frac{i\omega}{c^{2}}\vec{\bar{j}}-\nabla^{\prime}\bar{\rho}\right)G(z-z^{\prime},\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot})~,$ (170)
@@
$\vec{\widetilde{E}}_{\bot}(z,\vec{r}_{\bot})$ $=$ $-\frac{i\omega}{c}\bar{f}(\omega)\int_{0}^{z}dz^{\prime}\int d\vec{r^{\prime}}_{\bot}\exp\left\{i\omega\left[\frac{\mid\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}\mid^{2}}{2c(z-z^{\prime})}\right]+i\frac{\omega z^{\prime}}{2c\gamma^{2}}\right\}$ (182)
$\times\frac{1}{z-z^{\prime}}{\rho_{\bot}}\left(\vec{r^{\prime}}_{\bot}\right)\left(\frac{\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}}{z-z^{\prime}}\right)~.$
@@
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\times\int_{0}^{\infty}\frac{d\xi}{\xi}\exp\left[+i\omega\frac{\mid\vec{r}_{\bot}-\vec{r^{\prime}}_{\bot}\mid^{2}}{2c\xi}-\frac{i\omega\xi}{2c\gamma^{2}}\right]\Bigg{\}}$
@@
$\int_{0}^{\infty}{d\xi}\exp\left[i\left(-\xi+{\alpha}/{\xi}\right)\right]/{\xi}=2K_{0}\left(2~\sqrt{\alpha}\right)~,$ (186)
@@
$\vec{\widetilde{E}}_{\bot}(z,\vec{r}_{\bot})=\frac{i\omega\bar{f}(\omega)}{c}\exp\left[\frac{i\omega z}{2c\gamma^{2}}\right]\int d\vec{r^{\prime}}_{\bot}{\rho_{\bot}}\left(\vec{r^{\prime}}_{\bot}\right)$ (187)